OpenStudy (anonymous):

How do i calculate the frquences of the dominat and recessive alleles using this for say for people who can roll their tongue verse people who can not. ( 24/30 for yes ) (6/30) for no PLUS BB Bb and bb using p^2+2pq+q^2=1 ? PLEASE HELP i need to know this by tomorrow

OpenStudy (blues):

Figure out what p and q equal. The q^2 corresponds to the proportion of people who can't roll their tongue, which is 6/30. q^2 = 6/30 = 0.2. q = sqrt(0.2) <-- sqrt(0.2) means square root of 0.2 q = 0.4472. Then substitute that into the initial equation and solve for p. p^2 + 2pq + q^2 = 1 p^2 + 2*0.4472*p + 0.4472 = 1 p^2 + 0.8944p = 0.5288 p = 0.4064 Technically you get a second answer because it is a quadratic equation but it's negative. Negative frequencies are impossible, so you just ignore it. So the frequency of the dominant allele p is 0.4064 and the frequency of the recessive allele is 0.4472.